# Noncooperative Equilibrium Selection via a Trading-based Auction

### v2

@@ 4.3 Properties of TACo
There are several properties of the TACo procedure (Algorithm 1).
@@ 4.3.1 Procedural rationality
TACo incentivizes agents to act in their self-interest by enabling them to trade and negotiate based on their individual valuations of the available options. As demonstrated in Example 2, each agent selects the option that maximizes its profit, naturally driving the auction process toward a consensus.
@@ 4.3.2 Privacy preservation
One of TACo’s core strengths is its ability to preserve the privacy of each agent’s valuations. The trading mechanism functions without requiring agents to disclose their private valuations for the available options ($\mathcal{C}$) or trading assets ($\mathbf{b}$).
@@ 4.3.3 Independence from direct communication
TACo operates under a minimal communication framework, where agents only need to broadcast their choices. This eliminates the need for direct, 1-on-1 communication, making the algorithm straightforward to implement in practical systems.
@@ 5 Termination Proof of TACo
The intuition behind the convergence of TACo lies in that, as soon as a cycle is detected, the trading unit $d$ is reduced by a fixed decrement factor $\gamma$. However, as $d$ decreases, the profit differences within the cycle shrink and must thus eventually fall below the threshold $\epsilon$ for all agents. At this point, the agents reach a consensus by definition, and TACo terminates.
@@
Figure 2: (Illustration of Lemma 1, $n=4$) This figure states that the row sum $S_{i}^{k}$ of the $i$-th row of the profit matrix $\mathcal{J}$ remains constant every $n$ steps. The fluctuations in the row sum are bounded by the difference $S_{i}^{\text{max}}-S_{i}^{\text{min}}=d(n-1)b_{i}$, where $S_{i}^{\text{max}}$ and $S_{i}^{\text{min}}$ denote the maximum and minimum row sums during the cycle. The diagram highlights the periodic nature of the row sum across steps and the incremental fluctuations $db_{i}$ within each step.
@@ 5.1 Existence of cycles in TACo
The row sum $S_{i}^{k}=\sum_{j=1}^{m}\mathcal{J}_{i,j}^{k}$, where $m$ is the number of choices, remains constant every $n$ steps, with fluctuations equal to:
@@
Since the row sums of the profit matrix $\mathcal{J}$ remain constant over every $n$ steps (Lemma 1) and since $\mathcal{J}$ is lower-bounded (Lemma 2), it directly follows that $\mathcal{J}$ is also upper-bounded. We state this consequence in the following corollary.
@@
Exploiting the proved upper and lower bounds on $\mathcal{J}$ derived so far, we can now formally prove that TACo always enters a cycle. This result is important because, by definition, a state of consensus where all agents have similar preferences – up to a level of tolerance $\epsilon$ – is also a cycle. Hence, the convergence of TACo will subsequently boil down to showing that the cycle it eventually falls into corresponds to a consensus.
@@ 5.3 Finite termination property of TACo
With the profit differences bounded proportionally to $d$ (Theorem 2), we now demonstrate that TACo satisfies the epsilon-termination condition, leading to an approximate agreement in a finite number of steps. Recall that each time a cycle is detected – which provably happens according to Theorem 1 – the algorithm reduces the trading unit $d$ by a fixed decrement factor $\gamma$. This reduction ensures that the profit differences between choices eventually fall below the threshold $\epsilon$, leading to termination.
@@
We also provide a practical bound on the worst-case number of steps required till termination. This bound accounts for the number of cycles needed to meet the epsilon-termination condition as well as the steps within each cycle, hence providing practical insights into the efficiency of TACo.
Theorem 4 (Termination bound for TACo).
The TACo algorithm terminates at most within the following number of steps:
${\left\lceil\log_{\gamma}\left(\frac{\epsilon}{(m+1)d_{0}(n-1)b_{\max}}\right)\right\rceil\cdot n\cdot\big{(}(m+1)(n-1)\big{)}^{nm}},$ (19)
@@
Theorem 4 implies that if $\epsilon$ is increased to $\frac{\epsilon}{\gamma}$, then the worst-case termination bound decreases by $n\cdot\big{(}(m+1)(n-1)\big{)}^{nm}$ steps.
@@ 6 Numerical Experiments
6.1 Waypoint merging game
We model a common noncooperative scenario in air traffic management, where $n$ aircraft converging on a waypoint must adjust their speeds to avoid collisions, as shown in Figure 1. Each aircraft $i$ in this waypoint merging game has an estimated time of arrival (ETA) $e_{i}$ and seeks to minimize deviations from it while ensuring safe separation. We introduce carbon emission credits as a tradable secondary asset.
Each aircraft (agent) $i$ selects an action $x_{i}$, representing its adjustment to the ETA. The adjusted arrival time for agent $i$ is then $u_{i}=e_{i}+x_{i}$. The cost structure for each agent is:
$\begin{array}[]{ll}\underset{x_{i}}{\mbox{minimize}}&c_{i}(x_{i})=k_{i}x_{i}^{2}\\ \mbox{subject to}&|u_{i}-u_{j}|\geq D,\enskip\forall j\in[n]\setminus i,\end{array}$ (20)
where $k_{i}>0$ represents the urgency factor, indicating sensitivity to arrival time adjustments, and $D\geq 0$ is the required separation distance.
This game Equation 20 admits multiple generalized Nash equilibria. The absolute value in the constraint introduces two conditions depending on the arrival sequence. Specifically, the sign of the terms inside the absolute value changes based on the order of $u_{i}$ and $u_{j}$, leading to two inverted inequalities. This leads to $n!$ distinct equilibria. Thus, coordination is essential for safe and efficient merging.
Another key challenge is the lack of direct 1-on-1 communication between aircraft. Instead, aircraft use the Automatic Dependent Surveillance-Broadcast (ADS-B)system, which broadcasts publicly available information such as positions, urgency factors ($k_{i}$), and ETAs ($e_{i}$). However, ADS-B does not support direct negotiation, making consensus on an equilibrium difficult despite agents being able to compute all equilibria.
In our scenario, each aircraft privately determines its valuation for emission credits and employs the TACo algorithm to iteratively negotiate arrival adjustments. By trading emission credits and broadcasting preferences via ADS-B, aircraft reach a consensus while acting in their own self-interest, despite the communication constraints imposed by ADS-B.
@@ 6.3 Evaluation method
The simulations were implemented in the Julia programming language. We used the ParametricMCPs.jl package [41] and the PATH solver for mixed complementarity problems [42] to compute the Nash equilibria of the game Equation 20 prior to running TACo.
@@ 6.4 Results
Figure 3: Experimental results comparing the performance of the TACo algorithm with baseline methods (Voting, Utilitarian, Egalitarian, and Random dictator) across three evaluation metrics: optimality gap, Gini index, and termination steps. TACo demonstrates the smallest median values for the optimality gap and Gini index, highlighting its effectiveness in achieving near-optimal and fair outcomes.
@@ 6.4.1 Optimality and fairness
As shown in Figure 3(a), the TACo algorithm exhibited a consistently lower optimality gap compared to the baseline algorithms, except for Utilitarian, which achieves a zero optimality gap by definition. TACo showed minimal deviation from the optimal solution, with a median and third quartile value of 0% and a maximum value of 20.3%. This confirms that the TACo effectively minimizes deviations from the socially optimal solution.
Figure 3(b) shows that TACo achieves a low Gini index, with a median of 0.181, demonstrating that the TACo improves fairness by redistributing costs among agents. Notably, trading-based coordination enhances fairness beyond merely selecting the most balanced equilibrium. However, in cases where the cost differences between choices were initially small, the trading process amplified disparities, resulting in occasional outliers. This suggests that the choice of trading unit $d$ should account for the relative magnitude of costs in the system.
@@ 6.5.1 Effect of decrement factor on convergence
A similar trend was observed in the Gini index and optimality gap, as shown in Figure 5(b). Just as there was no significant improvement in the convergence rate for smaller $\gamma$ values, there was also no notable impact on optimality or fairness. Thus, as long as $\gamma$ does not exceed 0.9, there is no significant impact on convergence steps, optimality, or fairness.
@@ 7 Conclusion
Proof of Lemma 1.  Over every $n$ steps: 1. During its turn, agent $i$ updates $\mathcal{P}$ according to eq. 10, resulting in a payment of $d(n-1)b_{i}$ to other agents. 2. During the $n-1$ turns of other agents, agent $i$ receives $db_{i}$ from each, by updating the offer matrix $\mathcal{O}$ according to eq. 11.
@@
Note that in (28), the matrix $(n\textbf{I}_{n}-H_{n})$ is the Laplacian matrix of a fully connected graph with $n$ nodes. As shown in [43, p.8], for such a Laplacian one must have $\det(n\textbf{I}_{n}-H_{n})=0$, which indicates the existence of nontrivial solutions to (28). Transforming $(n\textbf{I}_{n}-H_{n})$ into row echelon form using Gaussian elimination yields:
@@
Since the row sum of the profit matrix is constant over every $n$ steps, this lower bound ensures the existence of an upper bound for the profits of the active choices. The upper bound, $U_{\sigma}^{i}$, occurs when one active choice achieves a profit value above the average, while the remaining $p-1$ active choices are at their lower bound. This situation leads to the following relation:
@@
Proof of Theorem 3.  By Theorem 1, the TACo process will always fall into a cycle. If the process enters a cycle involving a single choice, all agents consistently select the same choice. In this state, there is no profit difference, and the $\epsilon$-termination criterion is automatically satisfied.
For cycles involving multiple choices, from Theorem 2, the profit difference between any two choices for agent $i$ within a cycle is bounded as
@@
for all $j,j^{\prime}\in[m]$ and for all $i\in[n]$, where $k$ and $k^{\prime}$ are steps within the cycle. According to the trading unit reduction rule (item 4 in Section 4.2), when a cycle is detected (i.e., when a previously encountered profit matrix $\mathcal{J}$ reappears), the trading unit $d$ is reduced by a decrement factor $\gamma$, where $10<\gamma<$ 1. Each reduction in $d$ proportionally decreases the profit differences across choices, ensuring convergence toward zero.
@@
Since $\gamma<1$, and $m$, $n$, and $b_{\max}$ are constants, the term $(m+1)d_{0}(n-1)b_{\max}\cdot\gamma^{r}$ approaches zero as $r\to\infty$. Hence, for any $\epsilon>0$, there exists a finite $R$ such that for all $r\geq R$:
$\max_{(i,\mathcal{J}^{k}),(i,\mathcal{J}^{k^{\prime}})\in\sigma}|\mathcal{J}^{k}_{i,j}-\mathcal{J}^{k^{\prime}}_{i,j^{\prime}}|\leq\epsilon,\quad\forall i\in[n],\forall j,k\in A^{\sigma}.$ (44)
At this point, the $\epsilon$-termination criterion is satisfied, and the TACo process terminates.
Proof of Theorem 4.  The profit difference among the active choices for agent $i$ is given in Equation 39. Substituting this into Equation 25, we infer that the number of possible values that $\mathcal{J}_{i,j}$ can take in a cycle is at most $(p+1)(n-1)$, where $p$ is the number of active choices.
@@
$(m+1)d_{0}(n-1)b_{\max}\cdot\gamma^{r}\leq\epsilon,$ (45a)
$\textstyle{\implies r=\left\lceil\log_{\gamma}\left(\frac{\epsilon}{(m+1)d_{0}(n-1)b_{\max}}\right)\right\rceil,}$ (45b)
@@
${\left\lceil\log_{\gamma}\left(\frac{\epsilon}{(m+1)d_{0}(n-1)b_{\max}}\right)\right\rceil\cdot n\cdot\big{(}(m+1)(n-1)\big{)}^{nm}.}$ (46)
{IEEEbiography}
[[Uncaptioned image]]Jaehan Im (Graduate Student Member, IEEE) was born in Seoul, Republic of Korea. He received the B.S. and M.S. degrees in Aerospace Engineering from the Korea Advanced Institute of Science and Technology (KAIST). He is currently pursuing the Ph.D. degree in the Department of Aerospace Engineering and Engineering Mechanics at the University of Texas at Austin.
@@
[[Uncaptioned image]]David Fridovich-Keil (Member, IEEE) received the B.S.E. degree in electrical engineering from Princeton University, and the Ph.D. Degree from the University of California, Berkeley. He is an Assistant Professor in the Department of Aerospace Engineering and Engineering Mechanics at the University of Texas at Austin. Fridovich-Keil is the recipient of an NSF Graduate Research Fellowship and an NSF CAREER Award.

### v3

@@ 4.3 Properties of TACo
There are several properties of the TACo procedure.
@@ 4.3.1 Procedural rationality
TACo incentivizes agents to act in their self-interest by enabling them to trade and negotiate based on their individual valuations of the available options. This best-response decision rule corresponds to procedural rationality in the sense that agents make locally optimal choices under bounded information [39, 40]. As demonstrated in Example 2, each agent selects the option that maximizes its profit, naturally driving the auction process toward a consensus. Unlike voting-based methods, which may force majority outcomes that are not individually beneficial [41, 16], TACo ensures that the final outcome remains individually rational for all agents.
@@ 4.3.2 Privacy preservation
One of TACo’s core strengths is its ability to preserve the privacy of each agent’s valuations. The trading mechanism functions without requiring agents to disclose their private valuations for the available options ($\mathcal{C}$) or trading assets ($\mathbf{b}$), as only the agents’ current choice selections are broadcast at each round, and no cost or valuation information is revealed. This contrasts with bargaining-based schemes that rely on pairwise exchanges and require disclosure of utility or preference information [32, 16].
@@ 4.3.3 Independence from direct communication
TACo operates under a minimal communication framework, where agents only need to broadcast their choices. This eliminates the need for direct, 1-on-1 communication, making the algorithm straightforward to implement in practical systems. This stands in contrast to bargaining protocols that rely on repeated direct exchanges.
@@ 4.3.4 Termination guarantee
4.3.5 Consensus over discrete choices
TACo enables single discrete-choice consensus among noncooperative agents without central coordination. Conventional auction mechanisms, by contrast, are typically designed for resource allocation [35, 37, 22] and do not address this problem.
Remark 1 (Notes on rationality of TACo).
The update rule in TACo reflects procedural rationality under bounded and asymmetric information. Agents update their choices by selecting the option that maximizes their current profit at each iteration, which is the rational response given the information available to them at each iteration. Since reaching a joint decision is operationally essential in our setting, participation in the procedure is itself rational: abstaining would risk coordination failure, which is strictly more costly than any payment incurred through participation.
Remark 2 (Scope of rationality).
While the proposed update rule—iteration-wise best-response profit maximization—is not the uniquely rational behavior under all conceivable information structures, it is the justified strategy available under the privacy-preserving and decentralized conditions of TACo.
@@ 5 Termination Proof of TACo
The intuition behind the convergence of TACo lies in that, as soon as a cycle is detected, the trading unit $d$ is reduced by a fixed decrement factor $\gamma$. However, as $d$ decreases, the profit differences within the cycle shrink and must thus eventually fall below the threshold $\varepsilon$ for all agents. At this point, the agents reach a consensus by definition, and TACo terminates.
@@
Figure 2: (Illustration of Lemma 1, $n=4$) This figure illustrates that the row sum $S_{i}^{k}$ of the $i$-th row of the profit matrix $\mathcal{J}$ returns to the same value every $n$ steps. The fluctuations in the row sum are bounded by the difference $S_{i}^{\text{max}}-S_{i}^{\text{min}}=d(n-1)b_{i}$, where $S_{i}^{\text{max}}$ and $S_{i}^{\text{min}}$ denote the maximum and minimum row sums during the cycle. The diagram highlights the periodic nature of the row sum across steps and the incremental fluctuations $db_{i}$ within each step.
@@ 5.1 Existence of cycles in TACo
The row sum $S_{i}^{k}=\sum_{j=1}^{m}\mathcal{J}_{i,j}^{k}$, where $m$ is the number of choices, returns to the same value every $n$ steps, with fluctuations equal to:
@@
Since the row sums of the profit matrix $\mathcal{J}$ return to the same value every $n$ steps (Lemma 1) and since $\mathcal{J}$ is lower-bounded (Lemma 2), it directly follows that $\mathcal{J}$ is also upper-bounded. We state this consequence in the following corollary.
@@
Exploiting the proven upper and lower bounds on $\mathcal{J}$ derived so far, we can now formally prove that TACo always enters a cycle. By definition, consensus corresponds to the condition in which all agents have similar preferences—within a tolerance $\varepsilon$—which itself constitutes a cycle. Hence, the convergence of TACo will subsequently boil down to showing that the cycle it eventually falls into corresponds to a consensus.
@@ 5.3 Finite termination property of TACo
With the profit differences bounded proportionally to $d$ (Theorem 2), we now demonstrate that TACo satisfies the epsilon-termination condition, leading to an approximate agreement in a finite number of steps. Recall that each time a cycle is detected – which provably happens according to Theorem 1 – the algorithm reduces the trading unit $d$ by a fixed decrement factor $\gamma$. This reduction ensures that the profit differences between choices eventually fall below the threshold $\varepsilon$, leading to termination.
@@
We also provide a theoretical bound on the worst-case number of steps that TACo may spend within cycles until termination. This bound accounts for the number of cycles needed to meet the epsilon-termination condition as well as the steps within each cycle, hence providing practical insights into the efficiency of TACo.
Theorem 4 (Post-cycle termination bound for TACo).
The total number of steps that TACo can spend within cycles before termination is at most:
${\left\lceil\log_{\gamma}\!\left(\frac{\varepsilon}{(m+1)d_{0}(n-1)b_{\max}}\right)\right\rceil\cdot n\big((m+1)(n-1)\big)^{nm}},$ (14)
@@
Theorem 4 implies that increasing $\varepsilon$ to $\varepsilon/\gamma$ reduces the worst-case termination bound by one entire cycle, i.e., it decreases the maximum number of admissible cycles by one.
@@ 6 Numerical Experiments
6.1 Waypoint merging scenario
We model a common noncooperative scenario in air traffic management, where $n$ aircraft converging on a waypoint must adjust their speeds to avoid conflicts, as shown in Figure 1. Each aircraft $i$ in this waypoint merging scenario has an estimated time of arrival (ETA) $e_{i}$ and an urgency factor $k_{i}>0$ indicating its sensitivity to delay. The adjusted arrival time for aircraft $i$ is $u_{i}=e_{i}+x_{i}$, where $x_{i}$ denotes the chosen speed adjustment. We introduce carbon emission credits as a tradable secondary asset.
A central coordinator, such as a flow management unit, computes a set of safety-feasible arrival sequences. In our setting, we assume that the coordinator’s role is limited to ensuring safe separation and the selection among feasible options is left to the stakeholders. Formally, the coordinator minimizes the aggregate delay cost
$\begin{array}[]{ll}\underset{{\{x_{i}\}_{i=1}^{n}}}{\mbox{minimize}}&{\sum_{i=1}^{n}k_{i}x_{i}^{2},}\\ \mbox{subject to}&|u_{i}-u_{j}|\geq D,\enskip\forall i\neq j.\end{array}$ (15)
Because the sign of the terms inside the absolute value in Equation 15 changes depending on the arrival order, multiple feasible solutions can arise—up to $n!$ in total. The coordinator enumerates each solution representing a safe and system-feasible configuration, providing a set of options $\mathcal{X}=\{x^{(1)},x^{(2)},\ldots,x^{(n!)}\}$ to the aircraft.
Each aircraft privately evaluates these options according to its own urgency, fuel state, and valuation of carbon emission credits, resulting in heterogeneous preferences. Reaching agreement on a single shared plan therefore becomes a decentralized multi-choice consensus problem.
Another key challenge is the lack of direct one-on-one communication between aircraft. Instead, aircraft use the Automatic Dependent Surveillance-Broadcast (ADS-B) system, which broadcasts publicly available information such as positions and ETAs ($e_{i}$). However, ADS-B does not support direct negotiation, making it infeasible for aircraft to coordinate their preferences directly.
In this setting, each aircraft employs the TACo algorithm to reach agreement on one of the feasible options $\mathcal{X}$. By trading emission credits and broadcasting preference updates via ADS-B, aircraft achieve consensus while acting in their own self-interest.
@@ 6.3 Evaluation method
The simulations were implemented in the Julia programming language. We used the ParametricMCPs.jl package [43] and the PATH solver for mixed complementarity problems [44] to compute the solution of Equation 15 prior to running TACo.
@@ 6.4 Results
Figure 3: Experimental results comparing the performance of the TACo algorithm with baseline methods (Voting, Utilitarian, Egalitarian, and Random dictator) across three evaluation metrics: optimality gap and Gini index. TACo demonstrates the smallest median values for the optimality gap and Gini index, highlighting its effectiveness in achieving near-optimal and fair outcomes.
@@ 6.4.1 Optimality and fairness
As shown in Figure 3(a), the TACo algorithm exhibited a consistently lower optimality gap compared to the baseline algorithms, except for Utilitarian, which achieves a zero optimality gap by definition. TACo showed minimal deviation from the optimal solution, with a median and third quartile value of 0% and a maximum value of 20.3%. This empirically supports that the TACo effectively minimizes deviations from the socially optimal solution.
Figure 3(b) shows that TACo achieves a low Gini index, with a median of 0.181, demonstrating that the TACo improves fairness by redistributing costs among agents. Notably, trading-based coordination enhances fairness beyond merely selecting the most balanced choice. However, in cases where the cost differences between choices were initially small, the trading process amplified disparities, resulting in occasional outliers. This behavior arises when the intrinsic costs $\mathcal{C}$ are small relative to the trading unit $d$ and the termination tolerance $\varepsilon$, as the discrete trading updates can dominate the underlying cost structure and make outcomes appear less sensitive to the agents’ original cost profiles. This observation highlights the importance of selecting $d$ and $\varepsilon$ at scales consistent with the magnitude of $\mathcal{C}$, so that the intrinsic cost differences are properly reflected in the consensus outcome.
@@ 6.5.1 Effect of decrement factor on convergence
A similar trend was observed in the Gini index and optimality gap, as shown in Figure 5(b). Just as there was no significant improvement in the convergence rate for smaller $\gamma$ values, there was also no notable impact on optimality or fairness. Thus, as long as $\gamma$ does not exceed 0.9, there is no significant impact on convergence steps, optimality, or fairness. Intuitively, the decrement factor $\gamma$ controls the rate at which the trading unit $d$ shrinks, but it has limited influence on the eventual consensus outcome. A useful way to interpret this behavior is to imagine that the update dynamics effectively converge toward a region of agreement, with the $\varepsilon$–tolerance determining how close the trajectory must come before termination. Under this interpretation, $\gamma$ affects how quickly the process approaches this region, whereas $\varepsilon$ governs the final level of refinement. This explains why $\gamma$ influences convergence speed but shows little effect on fairness or optimality in our experiments.
@@ 6.5.2 Effect of algorithm interruption on convergence
6.6 Scalability test
Although Equation 14 provides a worst-case upper bound on the number of rounds required for termination, the bound is conservative. To evaluate the algorithm’s practical efficiency, we conducted a scalability test by varying both the number of agents ($n$) and the number of choices ($m$). We tested $n\in\{3,5,7,10\}$ and $m\in\{3,10,30,100\}$, using parameters $\gamma=0.9$, $\varepsilon=10^{-1}$, and initial trading unit $d_{0}=1$. For each $(n,m)$ pair, we sampled both the intrinsic cost matrix and the private valuations independently from uniform distributions, with $C_{ij},\,b_{i}\sim\mathrm{Uniform}(0,1)$. Each configuration was run for 100 trials, and we recorded the mean number of rounds to termination with confidence intervals.
As shown in Figure 7, the number of rounds grows sublinearly with respect to the number of choices $m$, whereas its dependence on the number of agents $n$ is significantly stronger, exhibiting growth that exceeds linear scaling for fixed $m$. These empirical observations indicate that TACo operates far more efficiently in practice than the theoretical worst-case bound Equation 14 suggests, while the bound remains useful for guaranteeing finite-time termination.
Figure 7: Scalability test of TACo, showing the number of rounds required until termination. For fixed number of agents ($n$), the rounds grow sublinearly with respect to the number of choices ($m$). For fixed $m$, the rounds increase at a superlinear rate as the number of agents ($n$) grows. Overall, the empirical trends indicate that TACo is far more efficient in practice than the theoretical worst-case bound in Equation 14.
@@ 7 Conclusion
Proof of Lemma 1.  Over every $n$ steps: 1. During its turn, agent $i$ updates $\mathcal{P}$ according to eq. 5, resulting in a payment of $d(n-1)b_{i}$ to other agents. 2. During the n1n-1 turns of other agents, agent $i$ receives $db_{i}$ from each, by updating the offer matrix $\mathcal{O}$ according to eq. 6.
@@
Note that in (23), the matrix $(n\textbf{I}_{n}-H_{n})$ is the Laplacian matrix of a fully connected graph with $n$ nodes. As shown in [45, p.8], for such a Laplacian one must have $\det(n\textbf{I}_{n}-H_{n})=0$, which indicates the existence of nontrivial solutions to (23). Transforming $(n\textbf{I}_{n}-H_{n})$ into row echelon form using Gaussian elimination yields:
@@
Since the row sum of the profit matrix is constant over every $n$ steps, this lower bound ensures the existence of an upper bound for the profits of the active choices. The upper bound, $U_{\sigma}^{i}$, occurs when one active choice achieves a profit value above the average, while the remaining p1p-1 active choices are at their lower bound. This situation leads to the following relation:
@@
Proof of Theorem 3.  By Theorem 1, the TACo process always enters a cycle. If the cycle involves a single choice, all agents repeatedly select the same choice; no profit differences arise, and the $\varepsilon$-termination condition is automatically satisfied.
For cycles involving multiple choices, Theorem 2 implies that the profit difference between any two choices for agent $i$ satisfies
@@
for all $j,j^{\prime}\in[m]$ and all $i\in[n]$. When a cycle is detected (i.e., when a previous profit matrix reappears), the trading unit $d$ is reduced by the factor $\gamma\in(0,1)$ (item 4).
@@
Since $\gamma<1$, and $m$, $n$, and $b_{\max}$ are constants, the term $(m+1)d_{0}(n-1)b_{\max}\cdot\gamma^{r}$ approaches zero as $r\to\infty$. Hence, for any $\varepsilon>0$, there exists a finite $R$ such that for all $r\geq R$:
$\max_{(i,\mathcal{J}^{k}),(i,\mathcal{J}^{k^{\prime}})\in\sigma}|\mathcal{J}^{k}_{i,j}-\mathcal{J}^{k^{\prime}}_{i,j^{\prime}}|\leq\varepsilon,\quad\forall i\in[n],\forall j,k\in A^{\sigma}.$ (38)
At this point, the $\varepsilon$-termination criterion is satisfied, and the TACo process terminates.
Proof of Theorem 4.  The profit difference among the active choices for agent $i$ is given in Equation 33. Substituting this into Equation 20, we infer that the number of possible values that $\mathcal{J}_{i,j}$ can take in a cycle is at most (p+1)(n1)(p+1)(n-1), where $p$ is the number of active choices.
@@
$(m+1)d_{0}(n-1)b_{\max}\cdot\gamma^{r}\leq\varepsilon,$ (39a)
$\textstyle{\implies r=\left\lceil\log_{\gamma}\left(\frac{\varepsilon}{(m+1)d_{0}(n-1)b_{\max}}\right)\right\rceil,}$ (39b)
@@
${\left\lceil\log_{\gamma}\left(\frac{\varepsilon}{(m+1)d_{0}(n-1)b_{\max}}\right)\right\rceil\cdot n\cdot\big((m+1)(n-1)\big)^{nm}.}$ (40)
{IEEEbiography}[[Uncaptioned image]]Jaehan Im (Graduate Student Member, IEEE) was born in Seoul, Republic of Korea. He received the B.S. and M.S. degrees in Aerospace Engineering from the Korea Advanced Institute of Science and Technology (KAIST). He is currently pursuing the Ph.D. degree in the Department of Aerospace Engineering and Engineering Mechanics at The University of Texas at Austin.
@@
[[Uncaptioned image]]David Fridovich-Keil (Senior Member, IEEE) received the B.S.E. degree in electrical engineering from Princeton University, and the Ph.D. Degree from the University of California, Berkeley. He is an Assistant Professor in the Department of Aerospace Engineering and Engineering Mechanics at the University of Texas at Austin. Fridovich-Keil is the recipient of an NSF Graduate Research Fellowship and an NSF CAREER Award.